A spring with a constant of 100N/m vibrates at 25Hz. What is the frequency of vibration of a spring with a constant of 400N/m?

1
Expert's answer
2021-12-09T15:48:20-0500

Given:

"k_1=100\:\rm N/m"

"f_1=25\:\rm Hz"

"k_2=400\:\rm N/m"

"f_2-?"


The frequency of vibration

"f=\frac{1}{2\pi}\sqrt{\frac{k}{m}}"

Hence


"f_2=f_1\sqrt{\frac{k_2}{k_1}}=2f_1=50\:\rm Hz"